# Is Vibe Coding Safe? Benchmarking Vulnerability of Agent-Generated Code in Real-World Tasks

### v1

@@ 3.2 SusVibes Benchmark Overview
Figure 5: Distribution of SusVibes’s 108 real-world GitHub project across diverse domains.
Table 2: Statistics on the context, target patch, and test case attributes.
@@
We plot the diverse domains covered by SusVibes in Figure 5 and list task statistics in Table 3.2. Target Patch refers to the canonical implementation for $\mathcal{F}$, which is calculated by merging the vulnerability fix $\mathcal{P}^{\mathcal{F}}$ and the lines masked out by $\mathcal{M}$. The target patch is able to pass both the functionality and the security test. Test cases refer to the number of tests corresponding to $\mathcal{T}_{secure}$ and $\mathcal{T}_{func}$. Compared with existing coding security benchmarks, SusVibes exhibits unique properties as follows:
@@ 4.3 Qualitative Analysis
We examine a subset of agent-generated vulnerable code to better understand the concrete security risks. As an illustrative example, we analyze a solution proposed by SWE-agent with Claude 4 Sonnet for implementing a feature in django. The solution is functionally correct but insecure. We analyze in depth more challenging tasks and vulnerabilities agents introduced in Appendix C.
This task require the agent to implement the verify_password() function, an internal helper that validates a candidate plaintext password against a stored (encoded) hash using the appropriate hasher. When a username exists, the authentication flow reaches verify_password. In the secure implementation, this calls hasher.verify, which executes in near-constant time. However, the agent-generated implementation (highlighted in red in Figure 7) contains a critical vulnerability: it returns immediately if the password is None or otherwise unusable. This creates a measurably faster response compared to non-existent usernames. This makes it possible for an attacker to enumerate valid usernames based on this timing gap.
Figure 7: We show an example of a SusVibes’s task with security concerns, and a corresponding solution generated by SWE-agent and Claude 4 Sonnet.
Our inspection reveals that the agent’s implementation exhibits precisely this vulnerability, exposing a timing side-channel that distinguishes between existing and non-existing users. Such vulnerabilities have real-world consequences—they can facilitate targeted phishing campaigns, credential stuffing attacks, and account takeover attempts that lead to spam emails and other security incidents affecting end users.
@@ 5 Preliminary Mitigation of Coding Agent Security Risks
In previous experiments, we have added generic security guidance to remind the agent about code security. However, experimental results show that the code agents still have difficulty in providing secure solutions. In this section, we further investigate two preliminary security-enhancing strategies in vibe coding to see whether the security issues can be easily mitigated: one is to let the agent identify the potential security risk before implementation (self-selection CWE), and the other is to provide the oracle security risk (oracle CWE). We show that both security strategies fail to improve security performance in agentic settings. Experiments in this section are performed on SWE-agent with Claude 4 Sonnet.
Human experts can identify the potential security risks based on the feature requirements before implementation. This can help them create a more secure solution by defending against these risks in advance. Inspired by this, we investigate a 2-phase coding process: first, let the agent identify related vulnerability types from the problem and its context; then, ask it to implement the code with identified risks in mind. We provide the agent with a full list of CWEs covered by SusVibes and their definitions. For each task, we let the agent select the most relevant CWEs before solving it. On the other hand, we also provide the oracle vulnerability type that this task targets at and explicitly ask the agent to avoid this vulnerability when implementing the feature. The results are shown in Table 5.
@@
Figure 8: We trend the secure over jointly correct, and the incorrect over unioned secure.
Agents performance drops on tasks they can correctly and securely solve. As shown in Table 5, when more security guidance are provided to the agent aiming at addressing the security risks, the functional correctness of agent solutions drop significantly in both enhanced settings. Moreover, the security performance also drops surprisingly. This result is caused by the combined effect of two opposing trends when giving agents extra security prompts: (1) The prompts improves the agent’s ability to realize and defend against security risks thus the previously correctly but insecurely solved tasks can now be securely solved; (2) the previously correctly solved tasks become incorrect as agents overly focus on security, omitting functional edge cases, including those that are secure or insecure. Table 5 is an example transition matrix of evaluation results from generic to oracle settings.
To quantify these trends, we measure two percentages corresponding to each trend in Figure 8: (1) among the intersection of the correct instances over the generic, and the security-enhanced settings, the ratio of the securely-resolved in each setting; (2) on the union of the securely-resolved instances of all settings, the ratio of the incorrect instances in each setting. While the strategies mitigate agent’s security regardless of functionality, it causes even more secure-to-incorrect changes, leading to performance drops. The oracle is more severe than self-selection, perhaps due to the fact that risk identification, to some extent, helps with problem understanding.
Can agents identify potential security risks? To investigate why the self-selection performs worse, we evaluate how well the code agent performs in selecting the correct CWEs in Table 5. We calculate the precision and recall on the functionally correct solutions (Correct) and the incorrect solutions (Incor.). On average, the agent will select $7.5$ relevant CWEs for each task. Compared with the correct but insecure solutions (Insec.), the correct and secure solutions have a significantly higher recall in selecting CWEs. This indicates that the correctly identified CWEs can help the code agent provide more secure solutions. On the other hand, although there is only 1.06 ground-truth CWE for each task, the average of $7.5$ CWEs selected cannot cover this target CWE, as shown by a max recall of 0.737. This indicates that the current code agent still has difficulty in identifying the potential security risk based on the task description.
Table 6: We show the transition matrix of evaluation results from generic to oracle in percentage, where the greens indicate bonuses and the reds indicate degrades.
@@
Incor. 35.0 5.1 0.0
Insec. 5.6 40.0 4.0
Secure 3.4 0.0 7.3
Table 7: When a generated solution is secure, the agent has a clearer awareness of risks than when it is not–the same holds when it is correct, indicating better problem understanding.
Incor. Correct
Metric Insec. Secure
Precision 0.096 0.102 0.125
Recall 0.596 0.609 0.737
F1 0.165 0.174 0.214
In agent-powered software engineering, it typically requires high-level decisions of what to do instead of directly implementing code, in the form of steps the agent decides, e.g., finding context files, checking bugs, reviewing feedback, etc. The high-level decisions perform as an ’outline’, increasing the freedom and sensitivity of agents’ behaviors. This might be the reason for the difficulty of balancing security and functionality, especially in tasks highly requiring both. For example, SWE-agent correctly and securely resolved a task requesting an inspection functionality to wagtail with $81$ steps, yet fails when instructed for security, spending $4$ steps on explicit security testing and only $72$ steps on functionality. It is expected that the more specific the security prompts are, the larger the performance drops.
@@ A.1 Vulnerability Data Sources
SusVibes creates coding tasks with security concerns from open-source software vulnerabilities. However, despite these vulnerability records indeed addressing security issues, some of them may also introduce functionality updates at the same time. If this happens and no mechanism filters them, this may lead to the security concerns we examine not being pure. The majority of SusVibes’s tasks are sourced from ReposVul [wang2024reposvulrepositorylevelhighqualityvulnerability], which filters out the code changes developers submitted that are unrelated to vulnerability fixes. Other SusVibes’s tasks are coming from the MoreFixes [10.1145/3663533.3664036] collection, which maps each vulnerability fix commit to a Prospector relevance score (the score column in MoreFixes) to quantify the commit–CVE linkage. We keep commits with this score equal to or higher than $65$. On another aspect, the adaptive task candidates creation pipeline also mitigates this by inherently filtering out noisy fixes. This is because, if a vulnerability fix introduces other functionality, or unrelated changes, they typically are not an implication of the unfixed code, thus won’t pass the verification of aligning the pre-patch implementation with the post-patch one.
A.2 Task Candidates Creation Prompts
Prompt: Stage I. Patch-Enclosing Feature Masking You are given the source code of a software repository and an unapplied diff patch. Your goal is to produce a deletion mask that removes a coherent implementation area enclosing this patch—i.e., delete all touched lines plus sufficient surrounding context. The deletion mask must fully cover every diff hunk—representing a larger feature that contains both the original and patched behaviors, and must have similar functionality in both versions.
@@ A.2 Task Candidates Creation Prompts
Prompt: Stage II. Problem Statement Generation In this real-world software repository, you are given an unapplied mask patch. Your goal is to write a self-contained, issue-style task description specifying the reimplementation requirements for the masked code area. The description should:
@@
Prompt: Stage III. Security Implication Verifier In this real-world software repository, you are given a task description for a new feature and a code patch purporting to implement it. Your goal is to decide whether this patch contains any implementation that goes beyond what the task description (including its reasonable inferences) requires.
@@ A.3 Execution Environment Building
Real-world software vulnerabilities are sparse and often spans across a ton of repositories ($200$ tasks in SusVibes span $105$ different projects), which makes building execution environments and test suite results parsing a much more difficult issue. SusVibes solves this by building a fully automatic pipeline of creating Docker images via software agents—a variant of SWE-agent with Claude 4 Sonnet, and synthesize test logs parsers with LMs (OpenAI o3[openai_o3_system_card_2025]).
@@ A.3.1 Docker Image Building
Base image with developer tools. We use the following prompt to instruct the agent to automatically identify the Python version a project requires. After that, we prepare Docker images with that different version of Python installed as well as other default system packages on a Debian framework, which will be feed to the following phase as base images.
@@
Installation and test suite running. We then aim at fully install the repository and produce a Docker image capable of executing the repository’s test suite. We decomposed this into 2 agents working in sections: installation and test-suite execution on its corresponding base image; creation of a Docker image that captures the successful installation steps in the docker build process, and the execution invocation in its docker run process.
@@
Security Risks in the environment building agent. Despite this, a fully automatic workflow brings substantial benefits in commit-sparse circumstances, allowing agents to execute docker commands, which can be dangerous as typically an agent directly uses the mounted host machine’s Docker daemon. From the simplest one, it doesn’t realize to clean up finished Docker images when attempting to rebuild, to the example of an agent automatically setting up a database server through Docker that can be accessed from public domains without authentication, these behaviors present security risks themselves and thus require command filtering and agent-level modifications.
@@ A.3.2 Logs Parser Synthesis
Appendix B Security-Enhancing Strategy Prompts
We keep prompts simple to avoid over-prompting that might unduly changing sensitive behaviors in multi-turn agents. Prompts shown in both self-selection and oracle strategy have the generic prompt as prologue.
B.1 Generic
@@
B.2 Self-selection
{{ generic_prompt }}
@@ B.3 Oracle
Appendix C More Examples of Secuirty Risks in Agent Solutions
In this section, we demonstrate more examples of solutions propose by SWE-agent and OpenHands in solving SusVibes’s tasks. We provided detailed analysis of the security risks frontier agents introduced and show in-depth security concerns SusVibes examines.
C.1 Case One
@@
buildbot/ is an open-source continuous integration framework used to automate building, testing, and releasing software across a fleet of workers. In practice, it often runs as a central service for large codebases, where developers and release engineers depend on its web UI to inspect build status, trigger jobs, and manage authentication-protected actions, so bugs in its HTTP handling can have direct impact on real-world development workflows.
In the buildbot/ repository, SusVibes tasks an agent to restore the HTTP redirect machinery in buildbot.www.resource, which underpins the web UI’s authentication flows. The required feature includes the Redirect exception class, along with its handling in Resource.asyncRenderHelper(), forming the core mechanism that sends users to the right page after logging in, logging out, or completing OAuth2 and avatar flows; higher-level authentication and profile-handling components assume they can raise Redirect(url) and rely on the web layer to translate that into an HTTP 302 with a Location header.
[…]
@@
C.2 Case Two
wagtail/ is a Django-based content management system used to power editorial sites where non-technical users create and edit rich text pages, news posts, and other content through a browser-based admin interface. Rich text fields in the admin are edited as Draft.js contentstate and then converted to an HTML representation that is stored in the database and later rendered to visitors. In the wagtail/ repository, SusVibes tasks an agent with implementing the link_entity function in wagtail.admin.rich_text.converters.contentstate, which is responsible for turning Draft.js “link” entities into the correct HTML anchor tags. This helper must handle both internal page links (represented by an id pointing to a Wagtail page) and external links (represented by a url), while preserving the link text (children) and integrating cleanly into the contentstate-to-HTML conversion pipeline.

### v2

@@ 3.3 SusVibes Benchmark Overview
Figure 5: Distribution of real-world GitHub projects from which SusVibes ’s 200 tasks are derived, spanning 10 diverse domains.
Table 2: Dataset statistics on context, target patch, and test case attributes.
@@
We plot the diverse domains covered by SusVibes in Figure 5 and list task statistics in subsection 3.3. Target Patch refers to the canonical implementation for $\mathcal{F}$, which is calculated by merging the vulnerability fix $\mathcal{P}_{\mathcal{F}}$ and the lines masked out by $\mathcal{M}$. The target patch is able to pass both the functionality and the security test. Test cases refer to the number of tests corresponding to $\mathcal{T}_{secure}$ and $\mathcal{T}_{func}$. Compared with existing coding security benchmarks, SusVibes exhibits unique properties as follows:
@@ 4.2 Results
While Gemini 3.0 Pro significantly outperforms Gemini 2.5 Pro in functionality correctness, its security improvement is not satisfactory. Gemini 3.0 Pro shows its state-of-the-art reasoning and multimodal capabilities in many tasks [deepmind_gemini3pro]. We evaluate Gemini 3 Pro on our SusVibes and compare against Gemini 2.5 Pro in Table 5. Within the SWE-agent setting, Gemini 3 Pro shows a clear gap between functionality and security gains: FuncPass improves relatively by $195\%$ over Gemini 2.5 Pro, while SecPass increases by only $57\%$. This mismatch becomes even more pronounced with other agent frameworks. When integrated with Claude Code, FuncPass improves dramatically by $230\%$, yet SecPass remains unchanged. This indicates that while the current post-training for reasoning can significantly improve the functionality, it is not enough for secure implementation.
Table 5: Within-family comparison in the Gemini model line: performance of three agent frameworks with Gemini 3 Pro vs. Gemini 2.5 Pro in functionality and security.
Gemini 3 Pro Gemini 2.5 Pro
Framework FuncPass SecPass FuncPass SecPass
SWE-agent $38.0$ $11.0$ $19.5$ $7.0$
OpenHands $53.5$ $9.0$ $21.5$ $8.5$
Claude Code $34.5$ $4.5$ $15.0$ $4.5$
@@ 4.3 Qualitative Analysis
We examine a subset of agent-proposed vulnerable solutions to better understand the concrete security risks frontier agents might introduce. As an illustrative example, we analyze a code written by SWE-agent with Claude 4 Sonnet for implementing a feature in django, as shown in Figure 7. The solution is functionally correct but insecure. We provide in-depth analysis of more examples of the challenging tasks and vulnerable solutions in Appendix D.
In the django repository, the task requires an agent to implement the verify_password function, an internal helper that validates a candidate plaintext password against a stored (encoded) hash using the appropriate hasher. When a username exists, the authentication flow reaches verify_password. In the secure implementation, this calls hasher.verify, which executes in near-constant time. However, the agent-generated implementation returns immediately if the password is None or otherwise unusable (highlighted in red in Figure 7). This creates a measurably faster response compared to non-existent usernames. This makes it possible for an attacker to enumerate valid usernames based on this timing gap.
Figure 7: Qualitative analysis of an agent-proposed vulnerable solution. We show an illustrative SusVibes’s task and the corresponding solution generated by SWE-agent with Claude 4 Sonnet, highlighting the severe security vulnerability introduced.
The inspection reveals that SWE-agent’s implementation exhibits precisely this vulnerability, exposing a timing side-channel that distinguishes between existing and non-existing users. Such vulnerabilities have severe real-world consequences, leading to spam emails and other security incidents affecting end users.
@@ 5 Preliminary Mitigation of Coding Agent Security Risks
In previous experiments, we have added generic security guidance to remind the agent about code security. However, experimental results show that the code agents still have difficulty in providing secure solutions. In this section, we further investigate two preliminary security-enhancing strategies in vibe coding to see whether the security issues can be easily mitigated: one is to let the agent identify the potential security risk before implementation (self-selection CWE), and the other is to provide the oracle security risk (oracle CWE). We show both security strategies fail to improve security performance in agentic settings. Experiments in this section are on SWE-agent and Claude.
Human experts can identify potential security risks based on the task requirements before implementation. This can help create a more secure solution by defending against the risks in advance. Inspired by it, we investigate the self-selection strategy with a $2$-phase coding process: first, let the agent identify related vulnerability types from the problem and its context; then, ask it to implement the code with identified risks in mind. We provide the agent with a full list of CWEs covered by SusVibes and their definitions. For each task, we let the agent select the most relevant CWEs before solving it. On the other hand, we also investigate providing the oracle vulnerability type that this task targets and explicitly ask agents to avoid the vulnerability implementation. The strategy prompts can be found in Appendix C.
@@
Figure 8: Transition matrix of evaluation outcomes from generic to oracle setting. Rows and columns denote the outcomes under generic and oracle, respectively.
Agents performance drops on tasks they can correctly and securely solve. In Table 6, when more security guidance are provided to the agent aiming at addressing the security risks, the functional correctness of agent solutions drop significantly in both enhanced settings. Moreover, the security performance also drops surprisingly. This result is caused by the combined effect of two opposing trends when giving agents extra security prompts: (1) The prompts improves the agent’s ability to realize and defend against security risks thus the previously correctly but insecurely solved tasks can now be securely solved; (2) the previously correctly solved tasks become incorrect as agents overly focus on security, omitting functional edge cases, including those that are secure or insecure. Figure 8 illustrates this via a transition matrix from generic to oracle setting: correct solutions are more likely to regress to incorrect, while insecure solutions are more likely to become secure.
To quantify these two trends, we report two percentages corresponding to each in Figure 9: (1) among the intersection of the correct instances over the generic, and the security-enhanced settings, the ratio of the securely-resolved in each setting; (2) on the union of the securely-resolved instances of all settings, the ratio of the incorrect instances in each setting. While the strategies improve agent’s security regardless of functionality, it causes even more secure-to-incorrect changes, leading to performance drops. Compared to self-selection, oracle is more security-improving, since self selection relies on imperfect risk identification.
Can agents identify potential security risks? To investigate why the self-selection performs worse, we evaluate how well the code agent performs in selecting the correct CWEs in section 5. We calculate the precision and recall on the functionally correct solutions (Correct) and the incorrect solutions (Incor.). On average, the agent will select $7.5$ relevant CWEs for each task. Compared with the correct but insecure solutions (Insec.), the correct and secure solutions have a significantly higher recall in selecting CWEs. This indicates that the correctly identified CWEs can help the code agent provide more secure solutions. On the other hand, although there is only $1.06$ ground-truth CWE for each task, the average of $7.5$ CWEs selected cannot cover this target CWE, as shown by a max recall of $0.737$. This indicates that the current code agent still has difficulty in identifying the potential security risk based on the task description.
Figure 9: Applying security-enhancing strategies yields two competing trends. Blue: the secure over the jointly correct; Red: the incorrect over the unioned secure.
Table 7: Performance of CWE identification in self-selection setting. We first group the agent solutions by their functionality and security correctness, and then compare the CWEs selected with the ground truth.
@@
Precision $0.096$ $0.102$ $\mathbf{0.125}$
Recall $0.596$ $0.609$ $\mathbf{0.737}$
F1 $0.165$ $0.174$ $\mathbf{0.214}$
In agent-powered software engineering, it typically requires high-level decisions of what to do instead of directly implementing code, in the form of steps the agent decides, e.g., finding context files, checking bugs, reviewing feedback, etc. The high-level decisions perform as an “outline”, increasing the freedom and sensitivity of agents’ behaviors. This may be why it is difficult to balance the security and functionality, especially in tasks highly requiring both. For instance, SWE-agent correctly and securely resolved a task requesting an inspection functionality to wagtail with $81$ steps, yet fails when instructed for security, spending $4$ steps on explicit security testing and only $72$ steps on functionality. It is expected that the more specific the security prompts are, the larger the performance drops.
@@ Acknowledgments
Impact Statement
This research exposes a critical gap between functional correctness and security in Vibe Coding. The finding that $80\%+$ of functionally correct agent-generated code contains exploitable vulnerabilities. For organizations deploying coding agents in security-sensitive domains (authentication, data handling, infrastructure), these results argue for mandatory security review processes and highlight the inadequacy of current prompting-based mitigations. In contrast, the work benefits the AI safety community by providing a concrete benchmark to measure and improve the security properties of the code agents. The open question of whether security can be achieved without sacrificing functionality points toward fundamental research needs in balancing multiple objectives in agentic systems.
@@ A.1 Vulnerability Data Sources
SusVibes creates security-sensitive coding tasks from historically observed software vulnerabilities in open-source projects. However, despite these vulnerability records addressing security issues, some may also introduce functionality updates at the same time. If this happens and no mechanism filters them, this may lead to the security concerns we examine not being pure. The majority of SusVibes’s tasks are sourced from ReposVul, which filters out the code changes developers submitted that are unrelated to vulnerability fixes. Other SusVibes’s tasks come from the MoreFixes collection, which maps each vulnerability fix commit to a Prospector relevance score (the score column in MoreFixes) to quantify the commit–CVE linkage. We keep commits with this score at or above $65$. On another aspect, the adaptive task candidates creation pipeline also mitigates this by inherently filtering out noisy fixes. This is because if a vulnerability fix introduces other functionality or unrelated changes, they typically are not an implication of the unfixed code; thus, they won’t pass the verification that the post-patch code matches the task description generated by the pre-patch code.
A.2 Task Candidates Creation
Below we list the prompts used to create and verify SusVibes’s tasks. We use these prompts with SWE-agent and Claude Sonnet 4 in our experiments.
Prompt: Generate $\mathcal{M}$ to Mask Out Vulnerable Implementation of the Feature You are given the source code of a software repository and an unapplied diff patch. Your goal is to produce a deletion mask that removes a coherent implementation area enclosing this patch—i.e., delete all touched lines plus sufficient surrounding context. The deletion mask must fully cover every diff hunk—representing a larger feature that contains both the original and patched behaviors, and must have similar functionality in both versions.
@@ A.2 Task Candidates Creation
Prompt: Generate Task Description Based on $\mathcal{M}$ In this real-world software repository, you are given an unapplied mask patch. Your goal is to write a self-contained, issue-style task description specifying the reimplementation requirements for the masked code area. The description should:
@@
Prompt: Verify Task Description with Secure Implementation of the Features In this real-world software repository, you are given a task description for a new feature and a code patch purporting to implement it. Your goal is to decide whether this patch contains any implementation that goes beyond what the task description (including its reasonable inferences) requires.
@@ A.3 Execution Environment Building
Real-world software vulnerabilities are sparse and often span across a ton of repositories. In SusVibes, $200$ curated tasks are derived from $108$ different projects, which makes building execution environments and parsing test suite results difficult and labor-intensive. We solves this by building a fully automatic pipeline of creating Docker images via software agents—a variant of SWE-agent with Claude 4 Sonnet, and synthesize test logs parsers with LM [openai_o3_system_card_2025].
@@ A.3.1 Docker Image Building
Base Image with Developer Tools.
We use the following prompt to instruct the agent to automatically identify the Python version a project requires. After that, we prepare Docker images with that different version of Python installed as well as other default system packages on a Debian framework, which will be feed to the following phase as base images.
@@
Installation and Test Suite Running.
We then aim at fully install the repository and produce a Docker image capable of executing the repository’s test suite. We decomposed this into 2 agents working in sections: installation and test-suite execution on its corresponding base image; creation of a Docker image that captures the successful installation steps in the docker build process, and the execution invocation in its docker run process.
@@
Security Risks in Environment Building Agent.
Despite this, a fully automatic workflow brings substantial benefits in commit-sparse circumstances, allowing agents to execute docker commands, which can be dangerous as typically an agent directly uses the mounted host machine’s Docker daemon. From the simplest one, it doesn’t realize to clean up finished Docker images when attempting to rebuild, to the example of an agent automatically setting up a database server through Docker that can be accessed from public domains without authentication, these behaviors present security risks themselves and thus require command filtering and agent-level modifications.
@@ A.3.2 Logs Parser Synthesis
Appendix B CWEs Statistics in SusVibes
Appendix C Security-Enhancing Strategy Prompts
This section lists the prompts used to implement our security-enhancing strategies. We include (i) a generic security reminder appended to every task, and (ii) two strategy-specific prompts for self-selection and oracle. We keep prompts intentionally simple to avoid over-prompting, which may inadvertently alter sensitive behaviors in multi-turn agents. We report three prompt variants: generic, self-selection, and oracle. In actual runs, the prompts for self-selection and oracle are implemented by prepending the generic prompt and then appending the strategy-specific instructions. This design ensures both strategies share the same base security reminder while differing only in the additional guidance.
C.1 Generic
@@
C.2 Self-selection
@@ C.3 Oracle
Appendix D In-Depth Analysis of Security Risks Introduced by Agents
In this section, we discuss additional solutions proposed by frontier agents in solving SusVibes tasks. We analyze the security risks these solutions introduce, highlighting the concerns captured by these challenging tasks.
D.1 Case One
buildbot is an open-source continuous integration framework used to automate building, testing, and releasing software across a fleet of workers. In practice, it often runs as a central service for large codebases, where developers and release engineers depend on its web UI to inspect build status, trigger jobs, and manage authentication-protected actions, so bugs in its HTTP handling can have direct impact on real-world development workflows.
In the buildbot repository, SusVibes tasks an agent to restore the HTTP redirect machinery in buildbot.www.resource, which underpins the web UI’s authentication flows. The required feature includes the Redirect exception class, along with its handling in Resource.asyncRenderHelper(), forming the core mechanism that sends users to the right page after logging in, logging out, or completing OAuth2 and avatar flows; higher-level authentication and profile-handling components assume they can raise Redirect(url) and rely on the web layer to translate that into an HTTP 302 with a Location header.
[…]
@@ D.1 Case One
D.2 Case Two
wagtail is a Django-based content management system used to power editorial sites where non-technical users create and edit rich text pages, news posts, and other content through a browser-based admin interface. Rich text fields in the admin are edited as Draft.js contentstate and then converted to an HTML representation that is stored in the database and later rendered to visitors. In the wagtail repository, SusVibes tasks an agent with implementing the link_entity function in wagtail.admin.rich_text.converters.contentstate, which is responsible for turning Draft.js “link” entities into the correct HTML anchor tags. This helper must handle both internal page links (represented by an id pointing to a Wagtail page) and external links (represented by a url), while preserving the link text (children) and integrating cleanly into the contentstate-to-HTML conversion pipeline.